OpenStudy (anonymous):

How do I simplify: (5x + 3/x + 2) + (6x - 1/ -2 - x)?

OpenStudy (eyust707):

shoot for that common deonominator

OpenStudy (anonymous):

5x + 3 6x -1 ------ + ------ x+2 -2 - x

OpenStudy (eyust707):

once we get that we are happy =)

OpenStudy (eyust707):

so how can we get both denominators to look a like

OpenStudy (anonymous):

(10x)+(3/x)+(5/2) would be the most simplified form

OpenStudy (anonymous):

I tried by I do this method: -2 - x = -+2 -- x = - (2-x) = ?

OpenStudy (eyust707):

factor out a -1

OpenStudy (eyust707):

or just multiply both top and bottom by -1

OpenStudy (anonymous):

so I multiply both top and bottom by -1 from the original problem?

OpenStudy (eyust707):

yes

OpenStudy (eyust707):

do it to the one on the right... itll make it easier

OpenStudy (anonymous):

ohhh..

OpenStudy (eyust707):

=)

OpenStudy (phi):

Did you get the answer?

OpenStudy (anonymous):

I got 5x + 3 -6x + 1 ------ + -------- x + 2 2+x

OpenStudy (anonymous):

but that's only part of it. eek!

OpenStudy (phi):

you know x+2 is the same as 2+x, so you have a common denominator. so now add the tops, and divide the whole thing by the common denominator

OpenStudy (anonymous):

ohhh.. it is? Then why is "x^2 - 4" and "4 - x^2" different?

OpenStudy (phi):

a+b = b+a (that's the x+2 = 2+x example) a-b does not equal b-a (that's the x^2 - 4 example)

OpenStudy (phi):

example 2+3= 3+2 (yes, 5) 2-3 = -1 3-2 = +1 not equal

OpenStudy (anonymous):

ohhh.. i see, i get it :P Thank you!

OpenStudy (phi):

what's the answer?

OpenStudy (anonymous):

- x + 4 ------- x+2

OpenStudy (phi):

From the book??

OpenStudy (anonymous):

both-- my answer and from the book.

OpenStudy (phi):

ok you could also write it as 4-x ---- x+2

OpenStudy (anonymous):

why?

OpenStudy (phi):

or \[ -\frac{x-4}{x+2} \]

OpenStudy (anonymous):

ohh..